OpenStudy (anonymous):

What is the equation of the line passing through (1, 7) and parallel to the line y = –x + 5 in standard form? Answer x – y = 8 x + y = 8 x + y = –8 x – y = –8

OpenStudy (anonymous):

\[y _{1}-y _{2}=m(x _{1}-x _{2})\] this is eq. of line

OpenStudy (anonymous):

y-7= m(x-1) i.e. filling in 1, and 7 for point. now to work out m (the slope)

OpenStudy (anonymous):

the slope of y=-x+5 is -1 because y=mx+c tells us m=-1

OpenStudy (anonymous):

our line is parallel to it so our equation becomes: y-7 = -1(x-1) y-7=-x+1 or y= -x + 1+7 => y= -x +8 now from our choices we see: y+x=+8 so the answer is y+x=8